OpenStudy (anonymous):

(6x-8)^1/2 = -2 solve for x?

OpenStudy (anonymous):

6x-8 = 4 6x= 12 x = 2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

if i was asked i would say this has no solution

OpenStudy (anonymous):

\[(6x-8)^{\frac{1}{2}}=-2\] \[\sqrt{6x-8}=-2\] and this means the positive number whose square is -2, and there is no such real number